This is from a worksheet: Sorry my equation editor is not letting me copy and paste the signs... hope this is not too hard to read. Thanks in advance!!
If sin 60 degrees = sq rt 3/2 and cos 60 degrees = 1/2, what are the values of
tan 60 degrees =
sec 60 degrees =
cot 60 degrees =
Then If csc (acute angle)= 3 and sec (acute angle) = 3 sqrt 2 over 4; what are the values of
sin (acute angle) =
tan (acute angle) =
cos squared (acute angle) =

Answer by Nate(3500)

If sin 60 degrees = sqrt(3/2) and cos 60 degrees = 1/2, what are the values of
tan 60 degrees = sin(60 degrees)/cos(60 degrees) = sqrt(3/2)/(1/2) = 2*sqrt(3/2)
sec 60 degrees = 1/cos(60 degrees) = 1/(1/2) = 2
cot 60 degrees = 1/tan(60 degrees) = 1/(2*sqrt(3/2))
Then If csc (acute angle)= 3 and sec (acute angle) = 3*sqrt(2)/4; what are the values of
acute angle: x
sin (acute angle) = csc(x)^-1 = 3^-1 = 1/3
tan (acute angle) = sin(x)/cos(x) = (1/cos(x))(1/sin(x))^-1 = (sec(x))(csc(x))^-1 = (3*sqrt(2)/4)(3)^-1 = sqrt(2)/4
cos squared (acute angle) = cos^2(x) = (cos(x))^2 = (1/cos(x))^-2 = (sec(x))^-2 = (3*sqrt(2)/4)^-2 = 4^2/(3*sqrt(2))^2 = 16/(9*2) = 8/9